OpenStudy (anonymous):

Which statement best describes how work and power are different? To find work we need to know force and distance; to find power we need to know force and velocity. To find work we need to know energy and time; to find power we need to know energy and distance. To find work we need to know velocity and distance; to find power we need to know distance and time. To find work we need to know distance and force; to find power we need to know energy and velocity.

OpenStudy (anonymous):

w=f*d p=w/t

OpenStudy (anonymous):

v=d/t

OpenStudy (anonymous):

use those to some how come up with formulas that satisfy 1 of your 4 descriptions